OpenStudy (anonymous):

Which equation is a quadratic equation? 3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3)

OpenStudy (anonymous):

the one so that when expanded yields a ax^2+bx+c=y form

OpenStudy (anonymous):

A isnt becaus subtracting 3x from both sides doesn't yield any x^2 term. --- D isn't because if you expand right hand side you get a 8x^3 term making it a cupic not a quadratic only leaves b and c and they kinda weird

OpenStudy (anonymous):

b and c seem to both be quadratic

OpenStudy (anonymous):

B for sure is so i'd go with b

OpenStudy (anonymous):

however if in doubt then lets expand C 7x2 + y = (3 + x)(7x − 1) y=(3 + x)(7x − 1)-7x^2 ^ ^ I can already tell we are going to have a 7x^2 term. And it will cancel with the -7x^2 term so it will be a line not a quadratic.